# 1pt plus 3pt minus 2pt

### v1

@@ Abstract
Human feedback plays a crucial role in transferring knowledge from humans to robots. Enhancing the expressiveness of human teaching is vital for both improving robots’ learning from humans and human-teaching-robot experience. In this work, we introduced a novel teaching signal: progress, designed to represent the completion percentage of a task. We conducted two online studies with 76 crowd-sourced participants and one public space study with 40 non-expert participants to validate the capability of this progress signal. For the public space study, we collected 40 demonstrations along with progress and scalar feedback through an ice cream topping-adding task. We find that progress could indicate whether the task is successfully performed, reflect the degree of task completion, identify unproductive but harmless behaviors, and may be more consistent across participants. Furthermore, our results show that giving progress does not require extra workload and time. An additional contribution of our work is a dataset of non-expert demonstrations from the public space study, which we observe to be multi-policy and sub-optimal, with sub-optimality not only from teleoperation errors but also from exploratory actions and attempts. The dataset is available at https://github.com/TeachingwithProgress/Non-Expert_Demonstrations [link redacted for anonymity].
@@ I Introduction
Robots have already firmly become part of our daily lives, making it crucial to learn from users, especially non-expert users. Learning from Demonstration (LfD) enables robots to learn new skills by revealing expert policies [brown2020better, chen2021learning] while Learning from Human Feedback (LfHF) allows robots to adapt to human preferences or correct wrong behaviors by learning or shaping a policy [cui2021empathic, cederborg2015policy, knox2009interactively]. More recent work has further shown that using human feedback and demonstrations together can make learning even more effective by reducing the data needs for human feedback [ibarz2018reward] and loosening the requirements of demonstrations to be near-optimal [brown2019ranking]. However, while interest in learning fully or partially from humans is high, there is relatively little research on what the most effective forms of human feedback are, especially in combination with human demonstrations.
Human feedback and human demonstrations can be complementary due to the difference in human knowledge they carry. Demonstrations carry relatively dense and global information including policies and goals, and tend to be less accurate [palan2019learning]. Feedback carries relatively sparse and local information such as the correctness or a rating of a robot’s action, and giving high-quality feedback can be much easier than giving high-quality demonstrations [arzate2020survey]. Perfect demonstrations are hard to obtain while purely learning from human feedback requires many human labels, often obtained at significant time and expense. To address these challenges and improve the quality of learning, human demonstrations can be combined with human feedback: demonstrations can be used to train an initial policy to improve the sample efficiency of feedback [ibarz2018reward], and feedback can be used to refine the policy learned from demonstrations [palan2019learning].
In previous work using human feedback alongside demonstrations, the forms of human feedback used are directly adapted from LfHF. The quality of a demonstration or a partial demonstration is typically assessed by comparing it to another [palan2019learning, brown2019ranking, brown2020better]. This approach of comparing or ranking demonstrations is extremely hard for naive users, especially when the trajectory is only a partial demonstration [laidlaw2021uncertain]. It also neglects the objective measurement of a demonstration or a partial demonstration itself and the fact that all demonstrations could be either good or bad. This is especially true for non-expert demonstrations: non-expert demonstrations can be noisy, multi-policy, and yet still succeed (we show this in subsection V-C). While feedback like binary evaluation and scalar feedback is capable of assessing the quality of a demonstration, comparative information is unavailable for binary feedback and is unreliable for scalar feedback [yu2023thumbs], especially when teachers are non-experts.
@@
In this work, we present a new type of human feedback: progress. Progress captures the degree of task completion. We show that progress can indicate the extent of task completion, determine if a task is completed, and be robust to unproductive behaviors. Furthermore, when compared to scalar feedback, progress is more consistent when demonstrations are noisy and does not need extra workload and time. We show the capability of progress through two online studies and one in-person study. For the online studies, we recruit 76 crowd workers to provide scalar feedback and progress with pre-recorded expert demonstrations over three simple tasks and one long-horizon task. For the in-person study, we recruit 40 passersby to provide demonstrations in an ice cream topping-adding task, shown in Figure 1. The demonstrators then provide progress and scalar feedback to their own demonstrations.
The main contribution of this work is the introduction of a new teaching signal progress that has information beyond rating and ranking, and has great potential in interactive learning. Our validation studies covered a wide range of scenarios, which involved two online studies, one in-person study, five different tasks, and 116 participants in total. Our results also showed that non-expert demonstrations are multi-policy and sup-optimal, but sub-optimal in a meaningful way. Finally, we released a dataset online with 40 non-expert labeled demonstrations from a public space study, which may better-reflect the types of demonstrations that can be expected from real-world deployments than typical expert or in-lab demonstration datasets.
@@ II Background
By leveraging human knowledge, interactive machine learning allows learning agents to adapt to the needs of individual users and improve sample efficiency [arzate2020survey]. Human knowledge can take a variety of forms, such as semantic representation [kostavelis2013learning], numerical feedback [knox2009interactively], eye gaze [saran2020understanding], gestures [yanik2013gesture], facial expressions [cui2021empathic], and demonstrations [ravichandar2020recent]. In this work, we focus on human feedback and human demonstrations. Separately, each of these approaches has significant limitations: on the one hand, inferring a policy from human feedback requires a large number of interactions [knox2009interactively, arakawa2018dqn, warnell2018deep], and on the other, error-free demonstrations are rare and expensive to obtain in the real world [ravichandar2020recent]. To address this, many approaches in Interactive Machine Learning seek to combine human feedback and human demonstration to compensate for the limitations of each.
@@
Learning from Human Feedback Learning from Human Feedback has emerged as a promising technology for robots or machine-learning agents to learn from humans via interactions [casper2023open]. LfHF, in general, refers to methods that have three components: feedback collection, policy or reward shaping, and policy optimization. Human feedback can be in a variety of forms, such as verbal [goyal2019using, kuhlmann2004guiding], numerical [cederborg2015policy, knox2009interactively, arumugam2019deep], and implicit [cui2021empathic, xu2021accelerating]. Three representative works of learning from explicit human feedback are Policy Shaping [cederborg2015policy], TAMER [knox2009interactively], and Preference-based policy learning [akrour2011preference]. Policy Shaping [cederborg2015policy] shapes the learned policy by querying users to indicate whether the robot’s action was correct. TAMER [knox2009interactively] replaces the environment rewards with human evaluations. Preference-based policy learning learns a policy based on preferences over two candidate policies. Human feedback has also been applied to modern Large Language Models (LLMs) [2023arXiv230308774O] to further improve the performance of trained models by having humans rate the outputs with binary critiques. Despite a large body of work has been done, the use of human feedback is mostly as reward signals. Thus, our work differs from prior work by proposing a new form of human feedback that carries not only rewards but also descriptions of task completion.
@@
Learning from Human Demonstrations In robotics, Learning from Demonstrations (LfD) is a method that facilitates robots to learn new skills by imitating humans [chernova2014robot]. The use of LfD offers several advantages, including eliminating the need for expert programming [zhu2018robot], high data efficiency [ravichandar2019skill], safety for learning [umlauft2017learning], and guaranteed performance [ijspeert2013dynamical]. The research interest in teaching robots via demonstrations has been steadily advancing. Even for complex continuous control tasks, LfD methods are capable of producing optimal behaviors with clean demonstrations and sufficient error-free demonstrations [sasaki2018sample]. However, due to the optimal assumption on the demonstrations, LfD methods like generative adversarial imitation learning [ho2016generative] or behavior cloning [fang2019survey] failed to acquire optimal policy for many robot tasks since even human experts would make mistakes while providing demonstrations [wu2019imitation]. One straightforward solution is screening out non-optimal demonstrations, but screening often involves a lot of human effort and significantly reduces sample efficiency because optimal demonstrations are rare [sasaki2020behavioral]. Instead of screening, weighting [sasaki2020behavioral, wu2019imitation, chen2021learning] or ranking [brown2020better, wang2021learning] demonstrations are considered to be more robust. However, learning from multiple users and learning from imperfect demonstrations are still challenging problems [ravichandar2020recent].
@@
Using Human Demonstrations and Human Feedback Recent work has demonstrated that combining human feedback and human demonstrations could overcome many disadvantages of using one of them solely, including safety [brown2020safe], sample efficiency [ibarz2018reward], and accuracy [palan2019learning]. Specifically, work from ibarz2018reward used demonstrations to train an initial model for efficiently collecting preferences from users. Work from palan2019learning built on [ibarz2018reward] and used a model-based method to reduce data from humans. Using human rankings, work from brown2019ranking has achieved supper human-demonstration performance. Although prior work has achieved great success in consolidating human feedback with human demonstrations, the source of human feedback and human demonstrations are likely from experts or pre-trained agents [huang2024modeling], and the human feedback they used could be more informative. Our work differs from prior work by focusing on human feedback and demonstrations from non-experts. We conducted our studies with crowd-sourced workers and random passersby. Our proposed new teaching signal is effective and consistent when human demonstrations are multi-policy and non-optimal.
@@ III Progress
Our goal is to improve learning from human feedback with a new teaching signal: progress. In this section, we first define progress, and then we introduce our hypothesis about progress.
@@ III-A Progress
We define progress as the accumulative task completion rate over a task based on the current observation, ranging from totally incomplete to fully complete. Our intuition is that: A teaching signal would be more robust to sub-optimal, and more consistent among users if human teachers could have objective references while providing the teaching signal. For progress specifically, users can use start states and finish states as references. In this work, we define the progress signal as a range from 0 to 100. A progress value of 0 indicates that the task has not yet begun, while a value of 100 signifies task completion. In between, we expect that for any given state $s$:
$prog(s_{i},a_{i},s)=prog(s_{j},a_{j},s)\forall i,j$ (1)
That is, progress is independent of the path taken to the state and is consistent, irrespective of the sequence of preceding states. However, human feedback is known to be noisy and can only be considered consistent if we view it as an approximate value [yu2023thumbs]. Thus, we collect progress by presenting users with a trajectory instead of a single state to increase reliability.
@@ III-B Hypotheses
We expect that people naturally estimate task progress in their daily lives, so progress should be not hard to give. We expect the increase and decrease in progress would indicate if the robot’s behavior is leading the task toward completion or away from completion. As a result, progress provided at the end of a trajectory should reflect if the task is completed. Thus, we hypothesize:
@@
H4.. Progress is more consistent than scalar feedback when demonstrations are non-optimal.
@@ IV Progress with Non-Self-Provided Demonstrations
We first recruited crowded users to provide progress with pre-recorded demonstrations. This allows us to explore the effectiveness of utilizing progress alone, and examine its applicability across a range of intricate scenarios, which we found common in our public space study. We conducted two online studies: Online study one uses three simple tasks to evaluate the workload of giving progress, while online study two involves a long-horizon task comprising six sub-tasks and five scenarios to assess the utility of progress.
@@ IV-A Online Study Setups
We believe that progress has information beyond linear time and scalar feedback. However, humans, especially non-experts, might find giving reasonable progress challenging resulting in requiring significantly more effort from users or little meaningful information in progress. To alleviate the concerns, we compared progress with scalar feedback in a range of 0 to 100. For each study, we recruited two groups of participants from an online platform. One group of participants was only asked to give progress and the other group of participants was only asked to give scalar feedback. Participants in two groups watched the same demonstrations in the same order.
@@
Figure 2: Online Study Setups. Three simple tasks for comparing workload of giving progress, and a long-horizon task for comparing applicability of progress.
IV-A1 Online Study One
@@
We recruited two groups of 20 participants to provide progress annotations to a robot performing three related tasks. The three tasks we used are moving, pouring, and spinning, shown in Figure 2, which are sub-tasks of the task we used in the online study phase two. For each task, participants gave 10 progress or 10 scalar feedback annotations. After giving all progress or scalar feedback, a NASA TLX [hart2006nasa] questionnaire was given to measure the workload. All demonstrations are perfect demonstrations except the robot made one mistake during the spinning task at step 5.
IV-A2 Online Study Two
@@
In study two, we used a long-horizon task with five representative types of variation in task performance to show the capability of progress. The task we use is a tea steeping task (also shown in Figure 2), which is a task combining picking, reaching, pouring, and spinning. We recruited another two groups of 18 participants from Prolific, a crowdsourcing platform for scientific research, to give scalar feedback or progress. We chose five demonstrations each representing a distinct scenario: perfect (everything is performed flawlessly), imperfect (the cup is dropped onto the table rather than being placed carefully), unaware (failed to pick the stir, and spun without a stir), corrected (failed to pick the stir but went back to pick it after a few spinning), and failure (the cup was dropped at the beginning). The results and further illustrations of five cases are shown in Figure 4. For demonstrations other than the corrected, participants were asked to give 15 progress or scalar feedback. In the corrected case, 20 progress or scalar feedback were collected.
IV-B Data Analysis
@@
To analyze the data, we used t-tests [witte2017statistics] and Bayesian statistics with the schemes present in [van2021jasp]. For t-test results, we will first use Shapiro-Wilk tests to determine if the data is not from a normal distribution. If the data is from a normal distribution, we will use a standard independent samples t-test. Otherwise, we will apply Kruskal-Wallis H Tests and Wilcoxon Rank-Sum Tests to our results. For Bayesian statistics, a Bayes Factor (BF) will be used. We interpret BF lower than 3 as ”no evidence” for the alternative hypothesis, between 3 to 10 as “moderate evidence”, and 30 or above as ”strong evidence”.
@@ IV-C Giving progress is not time-consuming and not hard
We used NASA TLX form to measure the workload, and we recorded the average task completion time. The results are shown in Figure 3. We did not find any significant difference between giving scalar feedback and progress in all the dimensions of the NASA TLX results using t-test and BF ($p>0.45$ and $BF<0.35$ for all dimensions). The average task completion time for evaluating the demonstrations using progress is 22 minutes 3 seconds, and the average task completion time for using scalar feedback is 23 minutes 10 seconds. The average completion time is lower for progress by $5.06\%$, but we did not find any significant difference using t-test and BF ($p<0.45$ and $BF<0.45$). H1 is supported.
Figure 3: Online Study One Results. The workload of giving progress has no difference from the workload of giving scalar feedback.
Figure 4: Online Study Phase 2 Results. Progress and scalar feedback are collected over five scenarios, Perfect, Imperfect, Unaware, Corrected, and Failure. Progress and scalar feedback carry different information. Scalar feedback reflects the optimality of a trajectory, while progress reflects the degree towards finishing the task successfully. Progress is more consistent than scalar feedback when the demonstration is non-optimal across participants. Progress is also capable of indicating if the task is completed successfully even if the robot has made a minor mistake or made a faulty mistake but fixed it.

### v2

# How Much Progress Did I Make? An Unexplored Human Feedback Signal for Teaching Robots

@@ Abstract
Enhancing the expressiveness of human teaching is vital for both improving robots’ learning from humans and the human-teaching-robot experience. In this work, we characterize and test a little-used teaching signal: progress, designed to represent the completion percentage of a task. We conducted two online studies with 76 crowd-sourced participants and one public space study with 40 non-expert participants to validate the capability of this progress signal. We find that progress indicates whether the task is successfully performed, reflects the degree of task completion, identifies unproductive but harmless behaviors, and is likely to be more consistent across participants. Furthermore, our results show that giving progress does not require extra workload and time. An additional contribution of our work is a dataset of 40 non-expert demonstrations from the public space study through an ice cream topping-adding task, which we observe to be multi-policy and sub-optimal, with sub-optimality not only from teleoperation errors but also from exploratory actions and attempts. The dataset is available at https://github.com/TeachingwithProgress/Non-Expert_Demonstrations.
@@ I Introduction
Robots have already firmly become part of our daily lives, making it crucial to learn from users, especially non-expert users. Learning from Demonstration (LfD) enables robots to learn new skills by observing expert policies [1, 2] while Learning from Human Feedback (LfHF) allows robots to adapt to human preferences or correct wrong behaviors by learning or shaping a policy [3, 4, 5]. More recent work has further shown that using human feedback and demonstrations together can make learning even more effective by reducing the data needs for human feedback [6] and loosening the requirements of demonstrations to be near-optimal [7]. However, while interest in learning fully or partially from humans is high, there is relatively little research on what the most effective forms of human feedback are, especially in combination with human demonstrations.
Human feedback and human demonstrations can be complementary due to the difference in human knowledge they carry. Demonstrations carry relatively dense and global information including policies and goals, and tend to be less accurate [8]. Human feedback carries relatively sparse and local information such as the correctness or a rating of a robot’s action, and giving high-quality feedback can be much easier than giving high-quality demonstrations [9]. Perfect demonstrations are hard to obtain while purely learning from human feedback requires many human labels, often obtained at significant time and expense. To address these challenges and improve the quality of learning, human demonstrations can be combined with human feedback: demonstrations can be used to train an initial policy to improve the sample efficiency of feedback [6], and feedback can be used to refine the policy learned from demonstrations [8].
In previous work using human feedback alongside demonstrations, the forms of human feedback used are directly adapted from LfHF [6, 8], which might not be effective for evaluating a demonstration. The quality of a demonstration or a partial demonstration is typically assessed by comparing it to another [8, 7, 1]. This approach of comparing or ranking demonstrations is extremely hard for naive users, especially when the trajectory is only a partial demonstration [10]. It also neglects the objective quality of a demonstration or a partial demonstration itself: a pair of demonstrations might both be good or both bad, making preferences difficult to provide. This is especially true for non-expert demonstrations: non-expert demonstrations can be noisy, multi-policy, and yet still succeed (we show this in subsection V-C). While feedback like binary evaluation and scalar feedback is capable of assessing the quality of a demonstration, comparative information is unavailable for binary feedback and is unreliable for scalar feedback [11].
@@
In this work, we characterize a novel type of human feedback for robot learning: progress, which is used to capture the completion of a task. We show that progress can indicate the extent of task completion, determine if a task is completed, and be robust to unproductive behaviors. Furthermore, when compared to scalar feedback, progress is more consistent when demonstrations are noisy and does not need extra workload and time. We show the capability of progress through two online studies and one in-person study. For the online studies, we recruit 76 crowd-sourced workers to provide scalar feedback and progress with pre-recorded expert demonstrations over three simple tasks and one long-horizon task. For the in-person study, we recruit 40 passersby to provide demonstrations in an ice cream topping-adding task, shown in Figure 1. The demonstrators then provide progress and scalar feedback to their own demonstrations.
The main contribution of this work is to demonstrate that progress has information beyond rating and ranking, and has great potential in interactive learning. Our validation studies covered a wide range of scenarios and involved 116 participants in total. Our results also showed that non-expert demonstrations are multi-policy and sub-optimal, but sub-optimal in a meaningful way. Finally, we released a dataset online with 40 non-expert labeled demonstrations from a public space study, which may better-reflect the types of demonstrations that can be expected from real-world deployments than typical expert or in-lab demonstration datasets.
@@ II Background
By leveraging human knowledge, interactive machine learning allows learning agents to adapt to the needs of individual users and improve sample efficiency [9]. Human knowledge can take a variety of forms, such as semantic representation [12], numerical feedback [5], eye gaze [13], gestures [14], facial expressions [3], and demonstrations [15]. In this work, we focus on human feedback and human demonstrations. Separately, each of these approaches has significant limitations: on the one hand, inferring a policy from human feedback requires a large number of interactions [5, 16, 17], and on the other, error-free demonstrations are rare and expensive to obtain in the real world [15]. To address this, many approaches in Interactive Machine Learning seek to combine human feedback and human demonstration to compensate for the limitations of each.
@@
Learning from Human Feedback Learning from Human Feedback has emerged as a promising technology for robots or machine-learning agents to learn from humans via interactions [18, 19] . LfHF, in general, refers to methods that have three components: feedback collection, policy or reward shaping, and policy optimization. Human feedback can be in a variety of forms, such as verbal [20, 21], numerical [4, 5, 22], and implicit [3, 23, 24]. Three representative works of learning from explicit human feedback are Policy Shaping [4], TAMER [5], and Preference-based policy learning [25]. Human feedback has also been applied to modern Large Language Models (LLMs) [26] to further improve the performance of trained models by having humans rate the outputs with binary critiques. Despite a large body of work that has been done, human feedback is mostly used as reward signals.
@@
Learning from Human Demonstrations In robotics, Learning from Demonstrations (LfD) is a method that facilitates robots to learn new skills by imitating humans [27]. The use of LfD offers several advantages, including eliminating the need for expert programming [28], high data efficiency [29], safety for learning [30], and guaranteed performance [31]. The research interest in teaching robots via demonstrations has steadily advanced. LfD methods are capable of producing optimal behaviors with clean demonstrations and sufficient error-free demonstrations [32]. However, due to the optimal assumption on the demonstrations, LfD methods like generative adversarial imitation learning [33] or behavior cloning [34] failed to acquire optimal policy for many robot tasks since even human experts would make mistakes while providing demonstrations [35]. Weighting [36, 35, 2] or ranking [1, 37] demonstrations are considered to be robust methods of learning from noisy demonstrations. However, learning from multiple users and learning from imperfect demonstrations are still challenging problems [15]. One previous work used a technique they refer to as “reward sketching”; in practice the annotators were instructed to provide progress [38]. While their work demonstrates the potential of progress for guiding learning, it did not closely investigate the properties of progress or take full advantage of it as a teaching signal, instead using large numbers of these “reward sketching” annotations as a loose approximation for a dense reward function.
@@
Using Human Demonstrations and Human Feedback Recent work has demonstrated that combining human feedback and human demonstrations could overcome many disadvantages of using one of them solely, including safety [39], sample efficiency [6], and accuracy [8]. Specifically, work from [6] used demonstrations to train an initial model for efficiently collecting preferences from users. Work from [8] built on [6] and used a model-based method to reduce data from humans. Using human rankings, work from [7] has achieved super-human demonstration performance. Although prior work has achieved great success in consolidating human feedback with human demonstrations, the source of human feedback and human demonstrations are likely from experts or pre-trained agents [40], and the human feedback they used could be more informative. Our work differs from prior work by focusing on non-experts. We conducted our studies with crowd-sourced workers and random passersby. We showed that progress is informative and consistent when human demonstrations are multi-policy and non-optimal.
@@ III Progress
Our goal is to improve learning from human feedback with a new teaching signal: progress. In this section, we first define progress, and then introduce our hypothesis.
@@ III-A Progress
We hypothesize that progress provides complementary information to demonstrations beyond rewards. We characterize progress as the accumulative task completion rate over a task based on the current observation, ranging from totally incomplete to complete fully. Our intuition is that: A teaching signal would be more robust to sub-optimal demonstrations and more consistent among users if human teachers could have objective references while providing the teaching signal. For progress specifically, users can use start states and finish states as references. In this work, we use the progress signal as a range from 0 to 100. A progress value of 0 indicates that the task has not yet begun, while a value of 100 signifies task completion. In between, we expect that for any given task $t$, current state $s$, any action $a_{i}$ and $a_{j}$, and any previous state $s_{i}$ and $s_{j}$:
$prog_{t}(s_{i},a_{i},s)=prog_{t}(s_{j},a_{j},s)\forall i,j$ (1)
Ideally, progress is independent of the path taken to the state, irrespective of the sequence of preceding states. However, human feedback is known to be noisy and can only be considered consistent if we view it as an approximate value [11]. Thus, we collect progress by presenting users with a trajectory instead of a single state to increase reliability.
@@ III-B Hypotheses
We expect that people naturally estimate task completion in their daily lives, so progress should be not hard to give. We also expect that participants would use progress to describe the completion degree of a task. We hypothesize:
@@
H4. Progress is more consistent than scalar feedback when demonstrations are non-optimal.
@@ IV Progress with non-self-provided demonstrations
We first crowd-sourced users to provide progress with pre-recorded demonstrations for our online studies. This allows us to explore the effectiveness of utilizing progress alone, and examine its applicability across a range of intricate scenarios. We conducted two online studies: Online study I uses three simple tasks to evaluate the workload of giving progress, while online study II involves a long-horizon task comprising six sub-tasks and five scenarios to assess the utility of progress.
@@ IV-A Online study setups
First, we validated the workload of providing progress and verified that progress contains unique information relative to scalar feedback in a range of 0 to 100. For each study, we recruited two groups of participants from an online platform. One group of participants was only asked to give progress and the other group of participants was only asked to give scalar feedback. Participants in two groups watched the same demonstrations in the same order.
@@
Figure 2: Online Study Setups. Three simple tasks for comparing the workload of giving progress, and a long-horizon task for comparing the applicability of progress.
IV-A1 Online study I
@@
We recruited two groups of 20 participants from Amazon Mechanical Turk to provide progress annotations to a robot performing three related tasks. The three tasks we used are reaching, pouring, and spinning, shown in Figure 2, which are sub-tasks of the task we used in the online study II. For each task, participants gave 10 progress or 10 scalar feedback annotations. After giving all progress or scalar feedback, a NASA TLX [41] questionnaire was given to measure the workload. All demonstrations are perfect demonstrations except the robot made one mistake during the spinning task at step 5.
IV-A2 Online study II
@@
In study II, we used a long-horizon task with five representative types of variation in task performance to show the capability of progress. The task we use is a tea steeping task (also shown in Figure 2), which is a task combining picking, reaching, pouring, and spinning. We recruited another two groups of 18 participants from Prolific, a crowdsourcing platform for scientific research, to give scalar feedback or progress. We chose five demonstrations each representing a distinct scenario: Perfect (everything is performed flawlessly), Imperfect (the cup is dropped onto the table rather than being placed carefully), Unaware (failed to pick the stir, and spun without a stir), Corrected (failed to pick the stir but went back to pick it after a few spinning), and Failure (the cup was dropped at the beginning). The results and further illustrations of five cases are shown in Figure 4. For scenarios other than Corrected, participants were asked to give 15 progress or scalar feedback. In Corrected, 20 progress or scalar feedback were collected.
IV-B Quantitative analysis
@@
To analyze the data, we used t-tests [42] and Bayesian statistics with the schemes present in [43]. For t-test results, we use Shapiro-Wilk tests to determine if the data is from a normal distribution. If the data is from a normal distribution, we use a standard independent samples t-test. Otherwise, we apply Kruskal-Wallis H Tests and Wilcoxon Rank-Sum Tests to our results. For Bayesian statistics, a Bayes Factor (BF) is used. We interpret BF lower than 3 as “no evidence” for the alternative hypothesis, between 3 to 10 as “moderate evidence”, and 30 or above as “strong evidence”.
@@ IV-C Giving progress is not time-consuming and not hard
We used the NASA TLX form to measure the workload, and we recorded the average task completion time. The results are shown in Figure 3. We did not find any significant difference between giving scalar feedback and progress in all the dimensions of the NASA TLX results using t-tests and BF ($p>0.45$ and $0.3<BF<1$ for all dimensions). The average task completion time for evaluating the demonstrations using progress is 22 minutes 3 seconds, and the average task completion time for using scalar feedback is 23 minutes 10 seconds. The average completion time is lower for progress by $5.06\%$, but we did not find any significant difference using t-test and BF ($p=0.442$, and $BF=0.444$). H1 is supported.
Figure 3: Online Study One Results. The workload of giving progress has no difference from giving scalar feedback.
Figure 4: Online Study II Results. Progress and scalar feedback carry different information. Scalar feedback reflects the optimality of a trajectory, while progress reflects the degree towards completing the task. Progress is more consistent than scalar feedback when the demonstration is non-optimal across participants. Progress is also capable of indicating if the task is completed successfully even if the robot has made a minor mistake or made a faulty mistake but fixed it.